OpenStudy (anonymous):

PLS HELP WILL PUT IN 4 PARTS FOR FOUR MEDALS IF YOUR NOT GONNA help dont say anything

OpenStudy (anonymous):

Bruce added some chlorine to the water in a pool. The chlorine evaporated at a fixed rate every week. The table below shows the amount of chlorine f(n), in ounces, that was left in the pool after n weeks: n f(n) 1 54 2 27 3 13.5 4 6.75 Which function best shows the relationship between n and f(n)? f(n) = 108(0.5)n−1 f(n) = 54(0.5)n f(n) = 108(0.5)n+1 f(n) = 54(0.5)n−1 Question 24 (Multiple Choice Worth 1 points) (04.06) A store had 175 cell phones in the month of January. Every month, 10% of the cell phones were sold and 10 new cell phones were stocked in the store. Which recursive function best represents the number of cell phones in the store f(n) after n months? f(n) = 175 − 0.9 × f(n − 1) + 10, f(0) = 175, n > 0 f(n) = 0.1 × f(n − 1) + 10, f(0) = 175, n > 0 f(n) = 175 + 0.9 × f(n − 1) + 10, f(0) = 175, n > 0 f(n) = 0.9 × f(n − 1) + 10, f(0) = 175, n > 0 Question 25 (Multiple Choice Worth 1 points) (04.03) Jack invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of his investment after x years: f(x) = 300(1.02)x What was the average rate of change of the value of Jack's investment from the third year to the fifth year? 6.43 dollars per year 8.24 dollars per year 12.86 dollars per year 14.26 dollars per year